OpenStudy (anonymous):

Suppose a section of the test has 50 multiple choice questions. Ron earns one point for each right answer and loses 1/4 point for each wrong answer. Ron answered 41 questions and his score was 34 3/4. How many questions did he answer correctly? how many questions did he answer incorrectly?

OpenStudy (anonymous):

MUST USE EQUATIONS

OpenStudy (anonymous):

can someone help me?

OpenStudy (anonymous):

Does he get any points if he does not answer?

OpenStudy (anonymous):

it does not include that information. so disregard that

OpenStudy (anonymous):

can you please help @purplez24

OpenStudy (sirm3d):

let x = no of items with wrong answer let y = no of items with correct answer equation 1: (total number of items answered) x + y = 41 equation 2: (points accumulated) 1y - (1/4)x = 34 3/4

OpenStudy (anonymous):

i agree the maximum possible points with all Q's answered would be 50 using only the info given and assuming nothing you would take 41 subtract his score then multiply the difference by four your answer will how many he got wrong

OpenStudy (anonymous):

I like to try with just 1 variable. 1x one point for each right answer. So x=right answers -1/4(41-x) we will lose 1/4 of a point for each wrong answer. (41-x) = wrong answers \[1x-\frac{ 1 }{ 4 }\left( 41-x \right)=34\frac{ 3 }{ 4 }\] \[x-\frac{ 41 }{ 4 }+\frac{ 1 }{ 4 }x =34\frac{ 3 }{ 4 }\] \[x+\frac{ 1 }{ 4 }x =34\frac{ 3 }{ 4 }+10\frac{ 1 }{ 4 }\] \[\frac{ 5 }{ 4 }x =45\] x=36 36 is the number of right answers. 41 - 36 = 5 is the number of wrong answers if we lose 1/4 of point for each wrong answer. then we will lose 1 1/4 points or 1.25 36 - 1 1/4 = 34 3/4. and that is what the problems says is our score